# Tackling Graph Oversquashing by Global and Local Non-Dissipativity

### v1

@@ 4.4 Ablation Study
In this section, we study the different components of SWAN, namely, the contribution of guaranteed global (i.e., graph-wise) and local (i.e., node-wise) non-dissipativity, and the benefit the spatial antisymmetry.
Table 3: Performance of different versions of SWAN on Graph Property Prediction and LRGB tasks. Results are averaged over 4 random weight initializations on the Graph Property Prediction, while over 3 on the LRGB. The first, second, and third best results for each task are color-coded.
Model Diam. SSSP Ecc. Peptides- Peptides-
func struct
$log_{10}$(MSE)$\downarrow$ $log_{10}$(MSE)$\downarrow$ $log_{10}$(MSE)$\downarrow$ AP $\uparrow$ MAE $\downarrow$
Weight Antisymmetry Only
SWANβ=0 -0.3882±0.0610 -3.2061±0.0416 0.5573±0.0247 0.6195±0.0067 0.2703±0.0023
Bounded Non-Dissipative
SWAN-ne -0.5497±0.0766 -3.1913±0.0762 0.3792±0.1514 0.6119±0.0037 0.2672±0.0012
SWAN-ne-learn -0.5631±0.0694 -3.5296±0.0831 0.1317±0.1253 0.6249±0.0051 0.2606±0.0007
Global and Local Non-Dissipative
SWAN -0.5249±0.0155 -3.2370±0.0834 0.4094±0.0764 0.6313±0.0046 0.2571±0.0018
SWAN-learn -0.5981±0.1145 -3.5425±0.0830 -0.0739±0.219 0.6751±0.0039 0.2485±0.0009
The importance of global and local non-dissipativity. To verify the contribution of the global and local non-dissipativity offered by SWAN, we also evaluate the performance of the following variants, which can deviate from being globally and locally non-dissipative, although in a bounded manner, as we discuss in Appendix E.1. Specifically, we consider the non-enforced (NE) variants of SWAN and SWAN-learn, which differ from the original version by the use of unconstrained weight matrix $\mathbf{V}$, rather than forcing it to be antisymmetric. These two additional variants are called SWAN-ne and SWAN-learn-ne, respectively.
Table 3 presents the performance of several SWAN variants on the Graph Property Prediction and LRGB tasks. The best performance on both synthetic and real-world problems is achieved through a global and local non-dissipative behavior by SWAN-learn, showcasing the importance of global and local non-dissipativity.
The benefit of spatial antisymmetry. Table 3 highlights the advantages of the spatial antisymmetry in Section 3. By setting $\beta=0$, we obtain a model with antisymmetry solely in the weight space, exhibiting only a local non-dissipative behavior. Our results showcase a noteworthy performance improvement when spatial antisymmetry is employed. This is further supported by the improved performance of -ne versions of SWAN, which do not guarantee both global and local non-dissipative behavior, compared to SWANβ=0. Thus, spatial antisymmetry emerges as a beneficial factor in effective information propagation.
@@ 5 Related Work
Graph Neural Networks based on Differential Equations. Adopting the interpretation of convolutional neural networks (CNNs) as discretization of ODEs and PDEs (Ruthotto & Haber, 2020; Chen et al., 2018b; Zhang et al., 2019) to GNNs, works like GCDE (Poli et al., 2019), GODE (Zhuang et al., 2020), GRAND (Chamberlain et al., 2021b), PDE-GCND (Eliasof et al., 2021), GRAND++ (Thorpe et al., 2022), DGC (Wang et al., 2021), and others, propose to view GNN layers as time steps in the integration of the heat equation. This perspective allows to control the diffusion (smoothing) in the network and understand the problem of oversmoothing (Nt & Maehara, 2019; Oono & Suzuki, 2020; Cai & Wang, 2020) in GNNs. Differently, Choromanski et al. (2022) propose an architecture with an attention mechanism based on the heat diffusion kernel. Other architectures like PDE-GCNM (Eliasof et al., 2021) and GraphCON (Rusch et al., 2022) propose to mix diffusion and oscillatory processes as a feature energy preservation mechanism. Other recent works have continued to explore this family of GNNs, by proposing additional mechanisms such as anti-symmetry (Gravina et al., 2023), reaction-diffusion-based dynamics (Wang et al., 2022; Choi et al., 2023), (Zhao et al., 2023; Eliasof et al., 2023) that propose models convection-diffusion and advection-reaction-diffusion to allow directed information transportation, and (Maskey et al., 2023) that formalize a fractional Laplacian ODE. While the aforementioned works, focus on the spatial aggregation term of DE-GNNs, as well as our SWAN, the temporal domain of DE-GNNs has also been studied recently in Eliasof et al. (2024); Kang et al. (2024). For a comprehensive review of differential equations and dynamical systems based GNN, we refer the interested reader to Han et al. (2023).
@@
Oversquashing in MPNNs. The issue of oversquashing in MPNNs has been widely recognized due to its impact on the inability of GNNs to effectively transfer information across nodes that are far apart in the graph (Alon & Yahav, 2021). In recent years, several strategies to alleviate oversquashing have been studied. One class of approaches involves graph rewiring, where methods such as SDRF (Topping et al., 2022) modify the original edge set to densify the graph as a preprocessing step. Similarly, GRAND (Chamberlain et al., 2021b), BLEND (Chamberlain et al., 2021a), and DRew (Gutteridge et al., 2023) dynamically adjust graph connectivity based on updated node features. Transformer-based approaches (Shi et al., 2021; Dwivedi & Bresson, 2021; Ying et al., 2021; Wu et al., 2023) opt to sidestep oversquashing by enabling message passing between all pairs of nodes. An alternative approach is the incorporation of non-local dynamics to construct dense graph communication functions, as seen in methods like FLODE (Maskey et al., 2023), which uses the fractional power of the graph shift operator. Similarly, QDC (Markovich, 2023) employs a dense quantum diffusion kernel, and G2TN (Toth et al., 2022) captures the diffusion path to mitigate oversquashing. Despite the success of these techniques in addressing oversquashing, a potential drawback is the increased complexity associated with propagating information at each update, often linked to denser graph shift operators. We also refer the reader to a recent survey in Shi et al. (2023).
@@
Non-dissipative Systems. Non-dissipative Systems are characterized by the absence of energy dissipation, thus playing a crucial role in various domains such as physics (Goldstein et al., 2002), engineering (Ogata, 2010), and machine learning. In the context of the latter, non-dissipative systems are utilized to model information flow efficiently. The effectiveness of a non-dissipative behavior for enhancing the computational power of reservoir computing and recurrent neural networks, has been demonstrated in Gallicchio et al. (2017) and Chang et al. (2019), respectively. In the context of GNNs, it was shown in Gravina et al. (2023) that local non-dissipativity is beneficial for long interaction modeling.
@@ A.1 Integration of SWAN
The ODE that defines SWAN follows the general DE-GNNs form, presented and discussed in Section 2.2. While there are various ways to integrate these equations (see for example various integration techniques in Ascher & Petzold (1998)), we follow the common forward Euler discretization approach, which is abundantly used in DE-GNNs literature (Chamberlain et al., 2021b; Eliasof et al., 2021; Gravina et al., 2023; Rusch et al., 2022). Formally, using the forward Euler method to discretize SWAN (Section 3.2) yields the following graph neural layer:
@@
Note, that in this procedure we replace the notion of time with layers, and therefore instead of using $t$ to denote time, we use $\ell$ to denote the step or layer number. Here, $\epsilon$ is the discretization time steps which replaces the infinitesimal $dt$ from Section 3.2. Furthermore, we note that similar to other DE-GNNs (Chamberlain et al., 2021b; Gravina et al., 2023) we use weight-sharing, such that the learnable matrices ${\bf W},{\bf V},{\bf Z}$, as well as the spatial aggregation matrices $\hat{{\bf A}},\tilde{{\bf A}}$ in $\Phi,\Psi$ (as shown in Equation 6, and discussed in A.2) are shared across layers.
Lastly, we notice that Euler’s forward method is stable if $(1+\epsilon\lambda(\mathbf{J}(t)))$ resides within the unit circle in the complex plane for all eigenvalues of the system (Ascher & Petzold, 1998). However, given that the Jacobian matrix’s eigenvalues are purely imaginary with our method, it implies $|1+\epsilon\lambda(\mathbf{J}(t))|>1$. Consequently, Equations (3.1) and (3.2) become unstable when solved using the forward Euler’s method. To strengthen the stability of the numerical discretization method, similar to Gravina et al. (2023) we introduce a small positive constant $\gamma>0$, which is subtracted from the diagonal elements of the weight matrix $\mathbf{W}$, with the aim of placing back $(1+\epsilon\lambda(\mathbf{J}(t)))$ within the unit circle.
@@ A.2 Spatial Aggregation Terms
In Equation 6 we utilize two aggregation terms, denoted by $\hat{{\bf A}}$ and $\tilde{{\bf A}}$. The first, $\tilde{{\bf A}}$, is used to populate the weight antisymmetric term $\Phi$, while the second, $\hat{{\bf A}}$, is used to populate the space antisymmetric $\Psi$. As discussed, in our experiments we consider two possible parameterizations of these terms, on which we now elaborate.
Pre-defined $\hat{{\bf A}},\ \tilde{{\bf A}}$. In this case we denote our architecture as SWAN and utilize the symmetric normalized adjacency matrix and the random walk normalized adjacency matrix for $\hat{{\bf A}},\ \tilde{{\bf A}}$, respectively. Formally:
@@
where ${\bf K}_{1}\in\mathbb{R}^{d\times 2d}$ and ${\bf K}_{2}\in\mathbb{R}^{n\times n}$ are learnable linear layer weights, and $\sigma$ is an activation function which is a hyperparameter of our method described in E.5. By averaging the feature dimension and gathering the averaged edge features $f_{(u,v)\in\cal{E}}^{emb}$ into a sparse matrix ${\bf F}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, such that ${\bf F}_{u,v}=\frac{1}{d}\sum f_{(u,v)\in\cal{E}}^{emb}$ we define the learned spatial aggregation terms as follows:
@@ A.3 The role of $\gamma$ and $\epsilon$
The $\gamma$ term in SWAN refers to a small positive constant subtracted from the diagonal elements of the weight matrix to balance the stability of the numerical discretization method (i.e., Euler’s forward method). According to Ascher & Petzold (1998), the Euler’s forward method applied to Section 3 is stable when $|1+\epsilon\lambda(\mathbf{J}(t))|>1$, with $\epsilon$ the step size. Since SWAN has $Re(\lambda(\mathbf{J}(t)))=0$, subtracting the constant $\gamma$ ensures that the Euler method remains stable. Therefore, as $\gamma$ approaches 0, the discretization method becomes increasingly unstable, requiring taking a smaller step size $\epsilon$.
@@ A.4 Applicability of SWAN to general MPNNs
Nowadays, most GNNs rely on the concepts introduced by the Message Passing Neural Network (MPNN) (Gilmer et al., 2017), which is a general framework based on the message passing paradigm. A general MPNN updates the representation for a node $u$ by using message and update functions. The first (message function) is responsible for defining the messages between nodes and their neighbors. On the other hand, the update function has the role of collecting (aggregating) messages and updating the node representation. Our SWAN, in Section 3, operates according to the MPNN paradigm, with the functions $\Phi$ and $\Psi$ that operate as the message function, while the sum operator among the node and neighborhood representations is the update function. Therefore, SWAN can be interpreted as a special case of an MPNN with the aformentioned parameterization, and can potentially be applied to different types of MPNNs.
@@
Let us assume a diffusion-based network whose Jacobian’s eigenvalues are represented by the diagonal matrix $\mathbf{\Lambda}$, and let denote the eigenvalues magnitude by a diagonal matrix ${\bf K}\in\mathbb{R}_{+}^{n\times n}$ such that ${\bf K}_{ii}=|\mathbf{\Lambda}_{ii}|$ for $i\in\{0,\ldots,n-1\}$. Applying the vectorization operator, it is true that :
@@
As it is known from Evans (1998), diffusion-based network are characterized by Jacobian’s eigenvalues with negative real part. Indeed, diffusion DE-GNNs are based on the heat equation. Therefore, the right-hand side of the ODE is the graph Laplacian, reading $\frac{\partial{\bf X}(t)}{\partial t}=-{\bf L}{\bf X}(t)$. Therefore, following our derivations in Section 2.3, we analyze $-{\bf L}$. It is known that the graph Laplacian has non-negative eigenvalues, and therefore $-{\bf L}$ has non-positive eigenvalues. Thus, the Jacobian has non-positive eigenvalues. Assuming a connected graph (i.e., it exists at least one edge), then there is at least one value that is strictly negative in $-{\bf L}$, and any entry that is not non-negative will be equal to zero. Therefore, we can write ${\bf J}=-{\bf K}$, leading to the equation
@@
Therefore, the information propagation rate among the graph nodes exponentially decays over time. Note, that at $t\rightarrow\infty$, it will converge to 1, which is exponentially lower than the propagation rate at early time $t$.
@@
This proof follows the one proposed in Di Giovanni et al. (2023) (Appendix B).
@@
We proceed by induction and we show only the inductive step (i.e., $\ell>1$), since the case $\ell=1$ is straightforward (we refer to Di Giovanni et al. (2023) for more details). Assuming the Einstein summation convention and given that $\mathbf{\hat{W}}=\mathbf{W}-\mathbf{W}^{\top}$, $\mathbf{\hat{Z}}=\mathbf{Z}+\mathbf{Z}^{\top}$, and $\alpha,\beta\in[p]$, we have:

### v2

@@ 4.4 Ablation Study
We now study the contribution of global (graph-wise) and local (node-wise) non-dissipativity, and spatial antisymmetry.
The importance of global and local non-dissipativity. To verify the contribution of the global and local non-dissipativity in SWAN, we evaluate the performance of several variants that can deviate from being globally and locally non-dissipative, although in a bounded manner, as discussed in Appendix E.1. Specifically, we consider the non-enforced (NE) variants of SWAN and SWAN-learn, which use an unconstrained weight matrix $\mathbf{V}$, rather than forcing it to be antisymmetric as in SWAN. These two additional variants are called SWAN-ne and SWAN-learn-ne, respectively. Table 3 shows the performance of the SWAN on the Property Prediction and LRGB tasks. The highest performance on synthetic and real-world problems was achieved with SWAN-learn, which is globally and locally non-dissipative.
The benefit of spatial antisymmetry. Table 3 highlights the advantages of the spatial antisymmetry in Section 3. By setting $\beta=0$, we obtain a model with antisymmetry solely in the weight space, exhibiting only a local non-dissipative behavior. Our results indicate a noteworthy performance improvement when spatial antisymmetry is employed. This is further supported by the improved performance of -ne versions of SWAN, which do not guarantee both global and local non-dissipative behavior, compared to SWANβ=0.
@@ 5 Related Work
Graph Neural Networks based on Differential Equations. Adopting the interpretation of convolutional neural networks (CNNs) as discretization of ODEs and PDEs (Ruthotto and Haber 2020; Chen et al. 2018) to GNNs, works like CGNN (Xhonneux, Qu, and Tang 2020), GCDE (Poli et al. 2019), GODE (Zhuang et al. 2020), GRAND (Chamberlain et al. 2021b), PDE-GCND (Eliasof, Haber, and Treister 2021), DGC (Wang et al. 2021), and others, propose interpreting GNN layers as discretization steps of the heat equation. This allows controlling the diffusion (smoothing) in the network and understanding the problem of oversmoothing (Nt and Maehara 2019; Oono and Suzuki 2020; Cai and Wang 2020) in GNNs. Differently, Choromanski et al. (2022) propose an architecture with an attention mechanism based on the heat diffusion kernel. Other architectures like PDE-GCNM (Eliasof, Haber, and Treister 2021) and GraphCON (Rusch et al. 2022) propose to mix diffusion and oscillatory processes as a feature energy preservation mechanism. Other recent works proposed mechanisms such as anti-symmetry (Gravina, Bacciu, and Gallicchio 2023), reaction-diffusion-based dynamics (Wang et al. 2022; Choi et al. 2023), and advection-reaction-diffusion (Eliasof, Haber, and Treister 2023). However, most of the aforementioned works are diffusion-based DE-GNNs, limited in modeling long-range interactions, while our SWAN is a DE-GNN with a constant information flow, addressing oversquashing in graphs. While the aforementioned works focus on the spatial aggregation term of DE-GNNs, the temporal domain of DE-GNNs has also been studied in (Eliasof et al. 2024; Gravina et al. 2024a; Kang et al. 2024; Gravina et al. 2024b). A review of these methods is given in Han et al. (2023).
@@
Oversquashing in MPNNs. Oversquashing in MPNNs, which hampers information transfer across distant nodes (Alon and Yahav 2021), has prompted various mitigation strategies. Graph rewiring methods like SDRF (Topping et al. 2022) densify graphs as a preprocessing step, while approaches such as GRAND (Chamberlain et al. 2021b), BLEND (Chamberlain et al. 2021a), and DRew (Gutteridge et al. 2023) dynamically adjust connectivity based on node features. Transformer-based models (Dwivedi and Bresson 2021; Rampášek et al. 2022) bypass oversquashing with all-to-all message passing. Another direction uses non-local dynamics to enable dense communication, as in FLODE (Maskey et al. 2023), which leverages fractional graph shifts, QDC (Markovich 2023) with quantum diffusion kernels, and G2TN (Toth et al. 2022), which captures diffusion paths. While effective, these methods often increase computational complexity due to dense propagation operators. For further discussion, see (Shi et al. 2023). We note that, long-range interactions have also been explored in sequential models (Hochreiter and Schmidhuber 1997; Gu, Goel, and Re 2022).
@@
Non-dissipative Systems. Non-dissipative Systems are characterized by the absence of energy dissipation, thus playing a crucial role in various domains such as physics (Goldstein, Poole, and Safko 2002), engineering (Ogata 2010), and machine learning, where such systems were shown to effectively model information flow. For example, the effectiveness of a non-dissipativity was demonstrated to enhance the power of reservoir computing (Gallicchio 2024) and recurrent neural networks (Chang et al. 2019). In the context of GNNs, it was shown in (Gravina, Bacciu, and Gallicchio 2023) that local non-dissipativity is beneficial for long interaction modeling.
@@ 6 Summary
Acknowledgments
The work has been partially supported by EU-EIC EMERGE (Grant No. 101070918). ME is funded by the Blavatnik-Cambridge fellowship, the Accelerate Programme for Scientific Discovery, and the Maths4DL EPSRC Programme.
@@ A.1 Integration of SWAN
The ODE that defines SWAN follows the general DE-GNNs form, presented and discussed in Section 2.2. While there are various ways to integrate these equations (see for example various integration techniques in (Ascher and Petzold 1998)), we follow the common forward Euler discretization approach, which is abundantly used in DE-GNNs literature (Chamberlain et al. 2021b; Eliasof, Haber, and Treister 2021; Gravina, Bacciu, and Gallicchio 2023; Rusch et al. 2022). Formally, using the forward Euler method to discretize SWAN (Section 3.2) yields the following graph neural layer:
@@
Note, that in this procedure we replace the notion of time with layers, and therefore instead of using $t$ to denote time, we use $\ell$ to denote the step or layer number. Here, $\epsilon$ is the discretization time step, which replaces the infinitesimal $dt$ from Section 3.2. Furthermore, we note that similar to other DE-GNNs (Chamberlain et al. 2021b; Gravina, Bacciu, and Gallicchio 2023) we use weight-sharing, such that the learnable matrices ${\bf W},{\bf V},{\bf Z}$, as well as the spatial aggregation matrices $\hat{{\bf A}},\tilde{{\bf A}}$ in $\Phi,\Psi$ (as shown in Equations 6 and 7, and discussed in Appendix A.2) are shared across layers.
Lastly, we notice that Euler’s forward method is stable if $(1+\epsilon\lambda(\mathbf{J}(t)))$ resides within the unit circle in the complex plane for all eigenvalues of the system (Ascher and Petzold 1998). However, given that the Jacobian matrix’s eigenvalues are purely imaginary with our method, it implies $|1+\epsilon\lambda(\mathbf{J}(t))|>1$. Consequently, Equations (3.1) and (3.2) become unstable when solved using the forward Euler’s method. To strengthen the stability of the numerical discretization method, similar to (Gravina, Bacciu, and Gallicchio 2023) we introduce a small positive constant $\gamma>0$, which is subtracted from the diagonal elements of the weight matrix $\mathbf{W}$, with the aim of placing back $(1+\epsilon\lambda(\mathbf{J}(t)))$ within the unit circle.
@@ A.2 Spatial Aggregation Terms
In Equations 6 and 7 we utilize two aggregation terms, denoted by $\hat{{\bf A}}$ and $\tilde{{\bf A}}$. The first, $\tilde{{\bf A}}$, is used to populate the weight antisymmetric term $\Phi$, while the second, $\hat{{\bf A}}$, is used to populate the space antisymmetric $\Psi$. As discussed, in our experiments we consider two possible parameterizations of these terms, on which we now elaborate.
Pre-defined $\hat{{\bf A}},\ \tilde{{\bf A}}$. In this case, we denote our architecture as SWAN and utilize the symmetric normalized adjacency matrix and the random walk normalized adjacency matrix for $\hat{{\bf A}},\ \tilde{{\bf A}}$, respectively. Formally:
@@
where ${\bf K}_{1}\in\mathbb{R}^{d\times 2d}$ and ${\bf K}_{2}\in\mathbb{R}^{n\times n}$ are learnable linear layer weights, and $\sigma$ is an activation function which is a hyperparameter of our method described in Appendix E.5. By averaging the feature dimension and gathering the averaged edge features $f_{(u,v)\in\cal{E}}^{emb}$ into a sparse matrix ${\bf F}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, such that ${\bf F}_{u,v}=\frac{1}{d}\sum f_{(u,v)\in\cal{E}}^{emb}$ we define the learned spatial aggregation terms as follows:
@@ A.3 The role of $\gamma$ and $\epsilon$
The $\gamma$ term in SWAN refers to a small positive constant subtracted from the diagonal elements of the weight matrix to balance the stability of the numerical discretization method (i.e., Euler’s forward method). According to (Ascher and Petzold 1998), the Euler’s forward method applied to Section 3 is stable when $|1+\epsilon\lambda(\mathbf{J}(t))|>1$, with $\epsilon$ the step size. Since SWAN has $Re(\lambda(\mathbf{J}(t)))=0$, subtracting the constant $\gamma$ ensures that the Euler method remains stable. Therefore, as $\gamma$ approaches 0, the discretization method becomes increasingly unstable, requiring taking a smaller step size $\epsilon$.
@@ A.4 Applicability of SWAN to general MPNNs
Nowadays, most GNNs rely on the concepts introduced by the Message Passing Neural Network (MPNN) (Gilmer et al. 2017), which is a general framework based on the message-passing paradigm. A general MPNN updates the representation for a node $u$ by using message and update functions. The first (message function) is responsible for defining the messages between nodes and their neighbors. On the other hand, the update function has the role of collecting (aggregating) messages and updating the node representation. Our SWAN, in Section 3, operates according to the MPNN paradigm, with the functions $\Phi$ and $\Psi$ that operate as the message function, while the sum operator among the node and neighborhood representations is the update function. Therefore, SWAN can be interpreted as a special case of an MPNN with the aforementioned parameterization and can potentially be applied to different types of MPNNs. Lastly, we note that, although we implemented $\Phi$ and $\Psi$ as Equations 6 and 7, their general formulation provides a flexible design principle for incorporating non-dissipativity as an inductive bias into any DE-GNN, while remaining accessible for non-experts users. Specifically, the proposed antisymmetric aggregation function ($\Psi$) can be easily implemented by leveraging ${\bf A}-{\bf A}^{\top}$ as the antisymmetric graph shift operator, while the weight matrix can be symmetrized using ${\bf Z}+{\bf Z}^{\top}$, making it compatible with standard methods such as GCN and GAT.
@@
Let us assume a diffusion-based network whose Jacobian’s eigenvalues are represented by the diagonal matrix $\mathbf{\Lambda}$, and let us denote the eigenvalues magnitude by a diagonal matrix ${\bf K}\in\mathbb{R}_{+}^{n\times n}$ such that ${\bf K}_{ii}=|\mathbf{\Lambda}_{ii}|$ for $i\in\{0,\ldots,n-1\}$. Applying the vectorization operator, it is true that :
@@
As it is known from (Evans 1998), diffusion-based networks are characterized by Jacobian’s eigenvalues with a negative real part. Indeed, diffusion DE-GNNs are based on the heat equation. Therefore, the right-hand side of the ODE is the graph Laplacian, reading $\frac{\partial{\bf X}(t)}{\partial t}=-{\bf L}{\bf X}(t)$. Therefore, following our derivations in Section 2.3, we analyze $-{\bf L}$. It is known that the graph Laplacian has non-negative eigenvalues, and therefore $-{\bf L}$ has non-positive eigenvalues. Thus, the Jacobian has non-positive eigenvalues. Assuming a connected graph (i.e., it exists at least one edge), then there is at least one value that is strictly negative in $-{\bf L}$, and any entry that is not non-negative will be equal to zero. Therefore, we can write ${\bf J}=-{\bf K}$, leading to the equation
@@
Therefore, the information propagation rate among the graph nodes exponentially decays over time. Note that at $t\rightarrow\infty$, it will converge to 1, which is exponentially lower than the propagation rate at early time $t$.
@@
This proof follows the one proposed in (Di Giovanni et al. 2023) (Appendix B).
@@
We proceed by induction and we show only the inductive step (i.e., $\ell>1$), since the case $\ell=1$ is straightforward (we refer to (Di Giovanni et al. 2023) for more details). Assuming the Einstein summation convention and given that $\mathbf{\hat{W}}=\mathbf{W}-\mathbf{W}^{\top}$, $\mathbf{\hat{Z}}=\mathbf{Z}+\mathbf{Z}^{\top}$, and $\alpha,\beta\in[p]$, we have: